the length of a rectangle is 8 cm more than its width. The perimeter is 68 cm. Find the length and width of the rectangle
1
Expert's answer
2012-12-03T09:09:18-0500
Length =x+8=L
Width=x=W
Perimeter=P=68cm
2W+2L=P
P=2x+2x+16=68
4x=68-16=52
X=13
So,& Length =13+8=21, Width=13
Answer: Length =21, Width=13;